# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalisation of Misinformation Detection Models

### v1

@@ 5.3 LLM Performance
    We compare the performance of the fine-tuned models to that of llama-3-8b-instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct Llama Team (2024), prompted to determine reliability of an article in a 0-shot setting (see App. E.1). We allow the model to ingest the first 512 tokens of the article [^7] Matching the amount of context the fine-tuned models receive.
    In Table 4 we compare the MCC score of the LLM to the 6 ‘Uniform’ fine-tuned models, evaluated on seen and unseen years. Despite the LLMs parameter count and the recency of its pre-training data, we find LLMs to be worse news reliability classifiers than the fine-tuned models. With the advent of LLM-based search engines, this is a somewhat worrying result, and aligns with recent results Liu et al. (2023); Yang and Menczer (2024).
    It should be noted that the use of even this modestly sized LLM incurs a computational cost far greater than fine-tuning. In our experiments, using a single A100 GPU, the LLM took ~$70$ hours to yield a prediction for all articles in the corpus, whereas each of the fine-tuning models were trained and evaluated on all remaining data in ~$12.5$ hours.
    Model MCC
    LLaMA-3 8B Instruct 0.25
    DeBERTa-v3 Unseen Year 0.39
    Seen Year 0.46
    Table 4: Comparing article-level MCC performance of an LLM classifier in a 0-shot setting to fine-tuned models.
    6 Analysis
@@ 6.1 Determinants of Performance
    The analysis of our results, thus far, has been constrained to article-level classification. While this reflects how misinformation classifiers interact with articles, it does not match how we annotate the dataset and can obscure performance on smaller, less mainstream publishers. Ideally, as highlighted by Baly et al. (2018b); Burdisso et al. (2024), classification performance is also evaluated at the publisher-level, testing which publisher properties aid or interfere with misinformation detection. To that end, we employ a binomial logistic regression on the average publisher-level accuracy to assess which aspects of a publisher determine the achieved accuracy score (for details and full model specification, see Appendix E.4). Unlike standard logistic regression, the dependent variable represents the ratio of correct classifications to all articles for each publisher individually (i.e., samples from a binomial distribution).
    In Figure 1 we show coefficient magnitudes for several important determinants, expressed as effect sizes Chinn (2000); Lampinen et al. (2022). A positive effect size indicates that the variable increases the odds of accurate classification, ceteris paribus. While interpreting the effect size magnitudes, the variable’s range of possible values must also be considered (see Appendix E.4 Table 16).
    The size of the training set has a large positive effect, with a $1.91$ multiplicative increase in the odds for each 10-fold increase in training samples. Thus, a publisher with 1000 articles in the training set is $3.82$ times more likely to have correct classification in the test set than a publisher with only 10 articles. Foreign publishers also prove easier to identify.
@@
    Figure 1: Coefficients of the determinants model, expressed as effect sizes. Circles are centred about the effect size, with lines giving the 95% confidence interval.
@@
    Relative to centre-biased publishers, publishers on the left or right sides of the political spectrum see slightly degraded performance, even after controlling for the publisher label. Moreover, classification on the unreliable-classes suffers more than on reliable+publishers. Since the odds ratios or effect sizes are difficult to interpret, we also provide the estimated marginal mean publisher-level accuracy for those combinations in Table 6.1.
@@
    Despite right biased unreliable-sources being far more prevalent in the training data (relative to left and centre bias), for both the ‘Questionable Source’ and ‘Conspiracy-Pseudoscience’ classes, their performance is worse. This somewhat confirms the results in the ‘Political Bias’ rows of Table 2: models struggle disproportionately to discriminate between reliable and unreliable-right-biased articles.
@@
    We see a slight positive correlation with the MBFC ‘Conspiracy-Pseudoscience’ level. The higher the value, the further the publishers’ articles tend to deviate from convention. As a result, strongly conspirational or pseudo-scientific publishers are $4.84$ and $4.72$ times more easily identified than publishers where this effect is weaker.
@@
    {tblr}
    colspec=rcccc, columns = colsep=2pt,
    Table 5: Predicted publisher-level accuracies averaged over combinations of the MBFC label (columns) and political bias (rows).
@@
    6.2 Effect of Publisher Diversity
    Finally, we test to what extent the diversity of the publishers present in the training data has an effect on the generalisation capacity of the models. We re-run the ‘Publisher’ split experiment with smaller training sets, constrained to only the most prolific publishers. Specifically, for each MBFC label, we only include the top-n most frequent publishers in the training data, while leaving the test set untouched. While this reduces the amount of variation in publishers considerably, it minimally affects the total amount of data present. Each training set still consists of hundreds of thousands of articles.
    Fig. 2 displays the generalisation gap (in terms of MCC) induced when increasing publisher homogeneity.
    Especially when limiting performance to the Top-1 most common publishers, the models show increased overfit to the training set. Where the ‘Publisher’ split saw a $0.1$ MCC delta, this increased to an average degradation of $0.5$. As the number of included publishers increases, the generalisation gap decreases and starts to converge to the previously seen ‘Publisher’ values. Notably, the variance in values is also substantially higher in the limited publisher settings.
    From this, we conclude that 1) the splitting described in Sec. 4 to have minimally altered the heterogeneity present in the dataset, and 2) the models improve with publisher heterogeneity.
    Figure 2: MCC scores for different ‘Publisher’ split test sets with only the top-1,2 or 3 most prolific publishers retained. The lightly shaded circles provide a value for each year of the dataset, the solid circles their average. The line in between provides the magnitude of the difference in performance across the seen and unseen publishers.
@@ 7 Conclusion
    This paper introduces misinfo-general, a dataset designed for training and evaluating out-of-distribution generalisation in misinformation detection models. We have defined and operationalised various axes of generalisation, some of which align with prior work, and some that are novel. We have shown that this dataset is challenging for models, and especially so when required to generalise to unseen forms of content or publishers.
@@
    Leveraging the metadata annotations available to us, we further analysed determinants of model performance, finding large discrepancies across political bias and misinformation forms, and have tentatively shown that increased diversity has a positive effect on generalisation.
@@
    While publisher-level labelling introduces noise, we believe the increased scale, diversity, and affluence of metadata make up for this. The first two properties can enable OoD generalisation or robustness, whereas the latter enables evaluation, analysis and potentially generalisation-aware training.
    We make the dataset publicly available, and hope it will serve as a resource for OoD generalisation focused training and evaluation. We hypothesise, but leave as an open research direction, that this large dataset might serve as a pre-training step to boost performance on smaller down-stream misinformation corpora. Otherwise, the use of publisher-level metadata might serve as a test-bed for generalisation focused misinformation detection algorithms.
    8 Limitations
@@
    While the dataset includes a diverse set of publishers, events and topics, ultimately the publisher metadata comes from a single source. The information provided by MBFC assumes a narrow, US-centric world-view. This is especially prominent when discussing foreign publishers. As such, the metadata provides limited nuance, providing only a single sample for an inherently subjective assessment. It is expected that across cultural backgrounds, the publisher information is bound to change.
@@
    In general, the news included in the dataset is US-centric, with the vast majority of publishers being American, producing articles for an American audience. This is exacerbated by us excluding all non-English articles. This means cross-lingual or cross-cultural generalisation cannot be evaluated.
@@
    That said, we do discuss one weak form of cross-cultural transfer. Besides prioritising different events, the primary distinction between the political ideologies discussed in Sec. 4, are their differing norms and values. As such, the poor political bias generalisation bodes ill for the more general cross-cultural generalisation tasks.
@@
    Specific to our approach, we experimented with only a single publisher to class label mapping. By binarising the labels, we leave no room for interpretation by the model. It is apparent that not all unreliable sources are equally unreliable, as shown in Sec. 6.1. Similarly, reliable sources can produce sensational content. Prior work has used different mappings, and we suggest future approaches to test various labelling schemes. With the provided metadata, we believe this can be easily implemented.
    9 Ethical Considerations

### v2

@@ 6.3 LLM Performance
We compare the performance of the fine-tuned models to that of llama-3-8b-instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct Llama Team (2024), prompted to determine reliability of an article in a 0-shot setting with $512$ token context (see Appendix D.2).
Despite the LLMs parameter count and the recency of its pre-training data, we find llama-3-8b-instruct to be inferior to the fine-tuned models as for the purpose of article-level reliability classification. It manages an MCC of $\mathbf{0.25}$, compared to our fine-tuning models achieving $\mathbf{0.46}$ on ID years, and $\mathbf{0.33}$ on OoD years. With the advent of LLM-based search engines, this is a somewhat worrying result, and aligns with recent findings of other works Liu, Zhang, and Liang (2023); Yang and Menczer (2024).
The use of this modestly sized LLM already incurs a computational cost far greater than the fine-tuning models used. In our experiments, using a single A100 GPU, the LLM took ~$70$ hours to yield a prediction for all articles in the corpus, whereas each of the fine-tuning models were trained and evaluated in ~$12.5$ hours.
7 Analysis of Model Generalization
@@ 7.1 Determinants of Performance
Figure 1: Coefficients of the determinants model, expressed as effect sizes. Circles are centered about the effect size, with lines giving the 95% confidence interval.
The analysis of our results, thus far, has been constrained to article-level classification. While this reflects how misinformation classifiers interact with articles, it does not match how we annotate the dataset and can obscure performance on smaller, less mainstream publishers. Ideally, as highlighted by Baly et al. (2018b) and Burdisso et al. (2024), classification performance is also evaluated at the publisher-level, testing which publisher properties aid or interfere with misinformation detection.
Table 4: Median predicted publisher-level accuracies averaged over combinations of the MBFC label (rows) and political bias (columns). The row headers correspond to (R) reliable, (Q) Questionable Source, (C) Conspiracy Pseudoscience, and (S) Satire.
{tblr}
[…]
Q 46.13% 46.11% 30.92%
C 32.47% 79.41% 31.20%
S 7.10% 14.11%
To that end, we employ a binomial logistic regression on the average publisher-level accuracy [^7] Defined as the ratio of correctly predicted articles to all articles for a publisher, i.e., true positive rate, recall. to assess which aspects of a publisher determine the achieved accuracy score (for details and full model specification, see Appendix E.3). Unlike standard logistic regression, the dependent variable is modelled as a ratio.
In Figure 1 we show coefficient magnitudes for several important determinants, expressed as effect sizes Chinn (2000); Lampinen et al. (2022). A positive effect size indicates that the variable increases the odds of accurate classification, ceteris paribus.
The size of the training set has a large positive effect, with a $\mathbf{1.91}$ multiplicative increase in the odds for each 10-fold increase in training samples. Thus, a publisher with 1000 articles in the training set is $\mathbf{3.82}$ times more likely to have correct classification in the test set than a publisher with only 10 articles. Foreign publishers also prove easier to identify.
@@
Relative to center-biased publishers, publishers on the left or right sides of the political spectrum see slightly degraded performance, even after controlling for the publisher label. Moreover, classification on the unreliable- classes suffers more than on reliable+ publishers. Since the odds ratios or effect sizes are difficult to interpret, we also provide the estimated marginal mean publisher-level accuracy for those combinations in Table 4.
@@
Despite right biased unreliable- sources being far more prevalent in the training data, for both the ‘Questionable Source’ and ‘Conspiracy-Pseudoscience’ classes, model performance is noticeably worse than on left and center biased sources. This somewhat confirms the results in the ‘Political Bias’ rows of Table 2: models struggle disproportionately to discriminate between reliable and unreliable- right-biased articles.
@@
We see a slight positive correlation with the MBFC ‘Conspiracy-Pseudoscience’ level. The higher the value, the further the publishers’ articles tend to deviate from convention. As a result, strongly conspirational or pseudo-scientific publishers are $\mathbf{4.84}$ and $\mathbf{4.72}$ times more easily identified than publishers where this effect is weaker.
@@
Figure 2: MCC scores for different ‘Publisher’ split test sets with only the top-1, 2 or 3 most prolific publishers retained per publisher class. The ‘Baseline’ column corresponds to the standard publisher splitting used in described in 5. The lightly shaded circles provide a value for each year of the dataset, the solid circles their average. The blue circles are ID publishers, the orange OoD publishers.
7.2 Effect of Publisher Diversity
Here we test to what extent the diversity of the publishers present in the training data has an effect on the generalization capacity of the models. We re-run the ‘Publisher’ split experiment with smaller training sets, constrained to only the most prolific publishers. Specifically, for each MBFC label, we only include the top-n most frequent publishers in the training data, while leaving the test set untouched. While this reduces the amount of variation in publishers considerably, it minimally affects the total amount of data present. Each training set still consists of hundreds of thousands of articles.
Figure 2 displays the generalization gap (in terms of MCC) induced when increasing publisher homogeneity. Especially when limiting performance to the Top-1 most common publishers, the models show increased overfit to the training set. Where the ‘Publisher’ split saw a $\mathbf{0.1}$ MCC delta, this increased to an average degradation of $\mathbf{0.5}$. As the number of included publishers increases, the generalization gap decreases and starts to converge to the previously seen ‘Publisher’ values. Notably, the variance in values is also substantially higher in the limited publisher settings.
From this, we conclude that (1) the splitting described in Section 5 to have minimally altered the heterogeneity present in the dataset, and (2) the models improve with publisher heterogeneity. The former finding suggests that the underestimation of the generalization gap will be especially egregious in datasets with a small pool of publishers (e.g., those that sample from a single reliable source to boost label balance). The latter, instead, provides some initial evidence for the utility of using large, diverse publisher-level datasets for pre-training article-level misinformation detectors; while fine-tuning on high-fidelity labels is likely necessary, using distantly supervised datasets might encourage more robust models before fine-tuning.
8 Analysis of Publisher-Level Labelling
8.1 Publisher Identifiability
The use of publisher-level labels as a form of weak supervision, especially in misinformation detection, can lead to models overfitting to publisher styles instead of article veracity. This was shown to be a serious concern by Zhou et al. (2021), and efforts to mitigate this effect at the data level were discussed in Section 4. Despite this, in Sections 6.1 and 7.1 we still found models to overfit to specific publishers and publisher classes, and in Section 7.2 we found a negative correlation between publisher diversity and the magnitude of the generalization gap.
@@ 8.1 Publisher Identifiability
Using the same learning algorithm, we find this to be a substantially more difficult task. While models exhibit above random article-level performance, with an average MCC score of $\mathbf{0.18}$, this is much lower than scores achieved with misinformation labels. Furthermore, when aggregating F1-scores across classes proportionally according to publisher frequency (micro) we get $\mathbf{0.14}$, whereas with a flat average (macro) we obtain a mere $\mathbf{0.04}$ F1. In short, while it is possible to predict the publisher from an article with above random performance, this is only really possible for the most prolific publishers, and this cannot entirely explain performance in misinformation classification.
8.2 Publisher Memorization
In this subsection, we analyze to which extent models need to memorize specific publishers. If there exists a lot of disagreement between the label of an article and the label assigned to its publisher, it is likely impossible to generalize to the unseen publisher from seen publishers; the labels of similar publishers clash. In this case, for classification to be successful, it is necessary for the misinformation detector to memorize publisher idiosyncrasies.
Inspired by the works of Pleiss et al. (2020), Jenkins, Talafha, and Goodwin (2023) in automated mislabelling detection, and Swayamdipta et al. (2020) on diagnosing dataset issues using ‘dataset cartography’, to estimate the necessity of memorization of specific publishers, we run an experiment comparing the average article confidence (mean logit assigned to the correct class) and disagreement (variance of logit assigned to the correct class) of publishers when included or excluded from the dataset. If there is a large shift in either the confidence or disagreement of a publisher when in- or excluded, this might indicate the publishers’ articles’ labels are not aligned with those of similar publishers.
We rerun the 2021 uniform split experiment 15 times. We exclude each publisher in 5 runs, at random, while taking care to minimize the number of exclusion set co-occurences and stratifying the exclusion across fine-grained MBFC publisher classes. As such, there should always be similar publishers available to excluded ones.
Figure 3 shows the effect of exclusion on the average confidence and disagreement scores for all publishers. Overall, and unsurprisingly, including a publisher during training increases average article confidence and decreases disagreement. However, for most publishers, the shift between training in- or exclusion is small, and likely attributable to the inherent stochasticity of mini-batch training. We assume that these publishers’ labels can largely be learned from similar publishers, and that these align well with each other.
Figure 3: The average article-level confidence and disagreement for different publishers. The top left panel shows scores when publishers are included in training, whereas the top right panel shows scores when publishers were excluded. The bottom left panel shows their difference, normalized. Only publishers with a shift magnitude above $2$ standard deviations (i.e., significantly far from origin) are shown with low opacity. The bottom right panel shows the direction of differences for significantly shifted publishers. In all panels, the size of the circles are proportional to a publisher’s size in the dataset. Green colored circles represent reliable+ publishers, whereas red colored circles represent unreliable- ones.
While there are significant shifts, these mostly occur for the largest, typically well-performing reliable+ , publishers, and take the form of significantly increased disagreement and decreased confidence when excluded. These include sources which MBFC believes to produce typically reliable, but sensationalist, subjective news (e.g., The Sun, The Daily Mirror), or anti-US propaganda sources (e.g., Pravda Report, Asia Pacific Research), as well as highly reputable sources (e.g., CBS News, BBC) [^8] These categorizations originate from MBFC, and do not reflect the authors’ opinions.. Comparatively, most large unreliable- publishers see far smaller shifts.
There are substantially fewer significant shifts in the other quadrants (Figure 3, bottom left panel), and those that do show up tend to be for much smaller publishers. Publishers that see a significant reduction in confidence when included in training do exist, although these comprise a small minority with typically very few articles. Looking more closely at such publishers, these include cases where site ownership changed during article collection (Viral News Network, Infinite Unknown), or whose articles are extremely noisy (Alternative Media TV), which would serve as good candidates for removal. We also find particularly difficult cases here, like neutral, objectively written articles promoting climate change denial (Climate Etc) [^9] Idem..
All in all, while there appears to be some ambiguity in article labels, largely due to publisher editorial biases, we find no evidence of mislabelling beyond a level expected for a corpus scraped from the internet. Despite having missed some publishers in the data cleaning phase (see Appendix B), these represent a small minority of all publishers, and collectively contain a small minority of all included articles.
8.3 Article Properties
In this subsection, we automatically analyze various properties of our dataset at the article level, both to assess their presence for different classes of publishers, and their correlation with news reliability labels.
Subjectivity analysis
The first property we annotate for is subjectivity. Objective news presents facts in a neutral, unbiased manner, and is commonly considered the antithesis of hyperpartisan or misinformation news, which is written specifically to incite an emotional response from readers, thereby inducing sharing Bojic, Prodanovic, and Samala (2024). Subjectivity has shown some promise as a feature in discriminating reliable and unreliable news Jeronimo et al. (2019). Despite this, reliable publishers also produce subjective text, likely to drive engagement. This can make articles unverifiable, hampering article-level labelling.
Figure 4: The estimated proportions for each subjectivity level in misinfo-general. Errors bars give an approximate 95% confidence interval Agresti and Coull (1998)
To assess the degree of objectivity, we ask ChatGPT-4o-mini to provide a rating for an article using a $5$-point Likert scale, ranging from entirely objective to entirely subjective. While by no means SoTA, similar setups have shown reasonable performance in prior work Galassi et al. ; Struß et al. ; Shokri et al. (2024).
Figure 4 shows the estimated proportions of each subjectivity level. Despite reliable news being in the majority, most ($\sim\mathbf{73}\%$) articles have a subjectivity level of ‘Mixed’ or higher. In fact, ‘Mostly Subjective’ articles seem to be most common.
Upon inspection of the dataset, these annotations seem to match our findings. While unreliable news is substantially less likely to present itself as objective, reliable publishers still publish a plethora of discussion and opinion pieces. This is especially true for publishers with a more pronounced political bias.
We repeat the binomial logistic regression analysis used in Section 7.1 to determine what publisher-level properties correlate with subjectivity. Unlike earlier, where the ease of classification correlated strongly with the form of misinformation, article subjectivity tends to correlate strongly with publisher political bias. Where an unreliable publisher reduces the odds of an objective article by between $\mathbf{0.34}-\mathbf{0.68}$, moving to a left or right political bias does so by between $\mathbf{0.24}-\mathbf{0.25}$. In other words, both reliable and unreliable publishers produce subjective, potentially unverifiable articles, especially when publishing from a biased standpoint. The full model, including estimated marginal medians, and prompt specification, can be found in Appendix E. We also compare the agreement with ChatGPT-4o, which seems to lean towards an even greater subjectivity propensity.
Figure 5: The association of emotions in articles with different publisher categories, as measured using pointwise mutual information. Each circle represents one of Plutchik’s 8 emotions, along with neutrality in their center. The area of a circle represents the strength of the association of the emotion with that publisher class, relative to the maximal association found. The legend provides each emotion’s color, moving clockwise from the top.
Emotion analysis
Another property we annotate for is emotion. Affective language in journalistic texts has long been understudied, despite emotion and its effect playing an increasingly important role in the modern media landscape Koivunen et al. (2021). It is especially prevalent in unreliable news, and is used to both persuade readers and incite sharing Alba-Juez and Mackenzie (2019). While the persuasiveness of emotional language in fake news is a matter of debate Martel, Pennycook, and Rand (2020); Phillips et al. (2024), with prior work showing that high affective state in people after ingesting misinformation being associated with both increased susceptibility Martel, Pennycook, and Rand (2020); Bago et al. (2022) and skepticism Horner et al. (2021); Lühring et al. (2024). From a computational perspective, however, combining emotion detection with misinformation detection has shown some promise Ghanem, Rosso, and Rangel (2020); Zhang et al. (2021b). Especially low valence emotions like anger, sadness, anxiety, surprise, and fear are believed to be prevalent in misinformation texts Liu et al. (2024).
We use a similar setup as above to annotate articles with one of eight Plutchik basic emotions Plutchik (1980) and neutrality, a common emotion model used for annotation in NLP Bostan and Klinger (2018).
Figure 5 shows the association of different emotions with different publisher classes. Visually, these results largely mirror the objectivity analysis; both reliable and unreliable publishers use emotional language, with political bias being a more important determinant of affect than reliability. The most notable difference in association is with ‘Neutral’ and Center-Reliable publishers; relative to other publisher categories, neutral writing occurs relatively often for this publisher class. Low valence emotions like ‘Anger’, ‘Disgust’ and ‘Sadness’ are prevalent throughout, but are especially associated with more politically biased, less reliable publishers. Especially ‘Satire’ publishers seem to be characterized by relatively high amounts of ‘Joy’ and ‘Surprise’. ‘Anticipation’ is highly prevalent in articles discussing expected future events, which appears to occur often, across all publisher classes.
Overall, however, much like subjectivity, there exists a good balance of emotions across the different publisher classes. While there is some correlation between emotion or subjectivity with publisher reliability, these appear to be insufficient to function as a shortcut for misinformation prediction.
While we partially corroborate the finding that different emotions are present (or at least, in different proportions) across publisher classes, ultimately there exists a good balance of emotions across the different publisher classes. Much like subjectivity, simply using the emotions present in an article to determine whether it comes from a reliable or unreliable publisher is likely insufficient.
As argued in Section 1, to reliably estimate the generalization gap of misinformation detectors, it is crucial to have access to a naturalistic corpus of misinformation, which is representative in terms of the diversity contained within. With this analysis, we have shown that for the properties of emotion and subjectivity, this diversity is present, with subjective and emotional language being present in articles from both reliable and unreliable publishers.
@@ 9 Conclusion
In the present state of the art, automated misinformation detectors cannot be safely or reliably deployed. While impressive performance is often reported, time and again, papers show that in more realistic settings, where out-of-distribution generalization is required, these models fail. We attribute this brittleness, backed by substantial evidence in prior work, to a failure of misinformation datasets to adequately simulate misinformation detection scenarios. Due to the prohibitively expensive cost of procuring misinformation labels, practically all surveyed misinformation datasets have had to make unrealistic assumptions, which introduce undesirable biases.
To accommodate recommended misinformation evaluation practices, and thereby enable the development of generalizable misinformation models, this article introduces misinfo-general. It consists of a benchmark built on top of a cleaned, weakly supervised corpus of online articles, which have rich article- and publisher-level metadata, and an operationalization of various generalization axes. We have shown that this dataset is challenging for a common class of misinformation detection models, and especially so when generalization to unseen forms of content or publishers is required.
@@
The metadata annotations enable us to further analyze the determinants of model performance. We find large discrepancies across political biases and misinformation forms, but have also shown that increased diversity has a positive effect on generalization.
@@
While publisher-level labelling introduces noise, we believe the increased scale, diversity, and affluence of metadata make up for this. The first two properties can enable OoD generalization or robustness, whereas the latter enables evaluation, analysis and potentially generalization-aware training.
We make the dataset publicly available, and hope it will serve as a resource for OoD generalization-focused training and evaluation. While the implementation of trustworthy automated misinformation detectors remains out of reach, we hope that this dataset at least makes evaluating and diagnosing generalization capacities of misinformation detection models easy enough for wide-spread academic adoption.
10 Limitations & Ethical Considerations
@@
While the dataset includes a diverse set of publishers, events and topics, ultimately the publisher metadata comes from a single source. The information provided by MBFC assumes a narrow, US-centric world-view. This is especially prominent when discussing foreign publishers from nations with geopolitical ambitions at odds with the U.S. As such, the metadata provides limited nuance, providing only a single perspective of an inherently subjective assessment. It is expected that across cultural backgrounds, the publisher information is bound to change.
@@
In general, the news included in the dataset is US-centric, with the vast majority of publishers being American, producing articles for an American audience. This is exacerbated by us excluding all non-English articles. This means cross-lingual or cross-cultural generalization cannot be evaluated.
@@
That said, we do discuss one weak form of cross-cultural transfer. Besides prioritizing different events, the primary distinction between the political ideologies discussed in Section 5, are their differing norms and values. As such, the poor political bias generalization bodes ill for the more general cross-cultural generalization tasks.
@@
Specific to our approach, we experimented with only a single publisher to class label mapping. By binarizing the labels, we leave no room for interpretation by the model. It is apparent that not all unreliable sources are equally unreliable, as shown in Section 7.1. Similarly, reliable sources can produce sensational content. Prior work has used different mappings, and we suggest future approaches to test various labelling schemes. With the provided metadata, we believe this can be easily implemented.